OpenStudy (anonymous):

What is the simplified form of (x^4 - 81)/(x + 3)?

OpenStudy (anonymous):

I found the common factors and I think it is x^3 – 3x^2 + 9x – 27.

OpenStudy (anonymous):

\[\frac{x^4 - 81}{x + 3}=\frac{(x^2 - 9)(x^2 +9)}{x + 3}=\frac{(x-3)(x+3)(x^2 +9)}{x + 3}\]Now,cancel the common factor, you'll get \[(x-3)(x^2 +9)\]If you expand it, you will get \[x^3-3x^2+9x-27 \]

OpenStudy (anonymous):

Okay cool, I got it then.